OpenStudy (dape):

I'm trying to find a proof that the congruence \[ x^2 \equiv a\mod m \] only has solutions for all integer \(a\) for \(m=1,2\).

OpenStudy (dape):

For example, when \(m=3\), there is no \(x\) such that \[ x^2\equiv2\mod 3.\] This is because any integer \(x\) can be written as either \(x_0=3k,\ x_1=3k+1\) or \(x_2=3k+2\). Squaring these three, we see that \[ x_0^2=9k^2 \\ x_1^2=9k^2+6k+1=3(3k^2+2)+1 \\ x_2^2=9k^2+12k+4=3(3k^2+6k+1)+1\] So the rest on division by 3 of \(x^2\) for any integer \(x\) is either 0 or 1. I'm pretty sure that there is always an \(a\) for which there is no solution when \(m>2\), but I haven't been able to prove it yet.

OpenStudy (kinggeorge):

Sorry for the late reply, but I just couldn't resist adding my input. Suppose \(m\ge 3\). Then \(x^2=1\) has two unique solutions: 1, and -1. So then if you take every element in the set \(S=\{0,1,...,m-1\}\) and square it modulo \(m\), then your resulting set of squares will have size at most \(m-1\lneq m\). So not every element in \(S\) can be a perfect square modulo \(m\). I think this will do it.

OpenStudy (dape):

Awesome, thanks! Really a case of over complicating things on my part.

OpenStudy (kinggeorge):

I've done my fair share of over complicating :P